# Line-of-sight effects on double source plane lenses

### v1

@@ 1 Introduction
Over the last few decades, a rich and varied array of cosmological probes have converged on what is now the standard model of cosmology: a spatially flat universe undergoing an accelerating expansion, dominated by cold dark matter and dark energy in the form of a cosmological constant [1, 2] Nonetheless, questions remain to be addressed before this model is universally accepted. Key tensions have emerged between early and late universe measurements of certain parameters, most notably $H_{0}$ and $\sigma_{8}$ [3, 4, 5, 6, 7]. Most probes suffer from astrophysical uncertainties, and these tensions limit the confidence with which other parameters can be constrained. Furthermore, tight constraints on dark energy dynamics have not yet been achieved, and recent results from the DESI collaboration seem to favour a redshift-dependent equation of state over the $\Lambda$CDM paradigm [8, 9].
@@
The first observation of such a system, The ‘Jackpot’ lens J0946+1006, was identified by [12] with the Sloan Digital Sky Survey [13] and follow ups with the Hubble Space Telescope. Later, a third, more distant multiply-imaged source was observed as part of the same system [14, 15], making J0946+1006 also the first triple source system. This system has been used to place constraints on $\Omega_{\mathrm{M}}$ and the dark energy equation of state parameter $w$. Most recently, re-analysis of the known galaxy-scale strong lensing system J1721+8842 revealed that the six observed quasar images arose from the same background source, the first confirmation of an Einstein zig-zag lens [16]. While only five other DSPL systems have been spectroscopically confirmed [17, 18, 19, 20, 21], surveys such as LSST and Euclid are expected to increase this number by at least two orders of magnitude [11].
In recent years, increasing attention has been paid to the impact on strong lensing observables of inhomogeneities along the line of sight, and it is now understood that, for precision applications of strong lensing, these effects cannot be ignored [22, 23, 24, 25, 26]. Gravitational shear effects arising from matter external to the main lens [27, 28, 29, 30, 31, 32] can be as strong as the shear produced by the lens itself [32], and a few percent of the total distortion caused by the main lens [33]. Consequently, azimuthal weak lensing effects must be included in models of strong lenses, and their omission can introduce biases [34, 35, 36] or prevent successful modelling [34]. The most common approach is to include a tidal “external shear” term, meant to absorb these effects, as well as any shear not reproduced by the mass model of the main lens [37, 38, 39, 32].
@@
This paper is organised as follows: in section 2, we review the theory of double source plane lensing and its use in constraining cosmology, and show how line-of-sight inhomogeneities affect DSPL observables. In section 3, we present our methods for simulating realistic tidal line of sight effects for DSPLs with arbitrary redshifts, and in section 4, we present the sample of DSPLs expected to be observed with Euclid. In section 5, we quantify the impact of the line of sight on the cosmological scaling parameter and the line of sight shear observed from DSPL systems. In section 6, we present flat-$w$CDM cosmological constraints from the Jackpot lens with an error budget updated to include a contribution from the line of sight, and then explore the impact of the line of sight on flat-$w$CDM and $w_{0}w_{a}$ cosmological constraints from the forecasted Euclid DSPLs. We summarise our results in section 7.
@@ 2.1.2 Cosmology with DSPLs
For a $w$CDM cosmology (with $w$ constant but not necessarily equal to $-1$), the normalised Hubble parameter $E(z)\equiv H(z)/H_{0}$ is given by
@@ 2.3.1 The cosmological scaling factor
where the $\sigma_{v}^{2}$ are the velocity dispersions of d and s1. Clearly, this relation need not be true in general.
@@ 4 Forecasted double source plane lens populations
Euclid is expected to contain a sample of $\sim 1700$ galaxy-scale DSPL systems that could plausibly be used to constrain cosmological parameters [72]. This forecast was derived from the LensPop package [73] modified to include multiple background sources and neglecting mass in the first source. This sample only includes systems where both sources have one or more arcs of length 0.30.3″0.3  ″, which ensures a reasonable possibility that the density slope of the lens can be recovered from Euclid imaging alone. In this work, we will assume that Euclid is able to find all of these forecasted lenses, with redshifts and Einstein radii following [72]. As in [72], we will assume that each DSPL enables a measurement of $\eta$ with an uncertainty given by the quadrature sum of 0.01 arcseconds over the Einstein radius of the first source, the same for the second source and 0.01. The first two terms represent reasonable uncertainties on the Einstein radius determined from lens modelling, whilst the last term sets a floor on the uncertainty given that the density profile slope of the primary lens is not perfectly known.
@@ 5.1 A single system - the Jackpot lens
From the figure on the left, we see that the line of sight typically introduces a bias on $\eta$ on the order of a percent, with a standard deviation of $\eta^{\mathrm{inferred}}/\eta$ of 0.67% for the inner pair of rings, and 1.3% for the outer. While the median of the bias is $\approx 0$, the histograms are not perfectly symmetric, and large overestimations of $\eta$ are more frequent than large underestimations. Unsurprisingly, we see that the typical size of the bias increases when considering the value of $\eta$ as calculated between s1 and s3, as the larger redshift differences between these two sources mean the line of sight becomes, on average, more significant (see fig. 1). For comparison, the measurement error in $\eta$ presented in [60] is 1.1%. That statistical error thus remains the dominant source of uncertainty, but the effect of the line of sight is not completely negligible.
From the figure on the right, we see that the difference between the line-of-sight shear acting on the inner and outer rings is typically slightly smaller than a percent, once again growing larger when comparing with the more distant source. While these values are small in absolute terms, this corresponds to a median difference of $\sim$140of $\eta\,g_{\mathrm{LOS,1}}$ in the case of s2, and $\sim$230in the case of s3. It is therefore clear that $\eta\,g_{\mathrm{LOS,1}}$ is a poor substitute when modelling $g_{\mathrm{LOS,2}}$ and $g_{\mathrm{LOS,3}}$. Compound lens models should allow for an additional external shear term on each lens plane.
@@ 5.2 A sample of mock Euclid DSPLs
While the field of double-source plane lensing has been previously restricted to a handful of systems, surveys with telescopes such as the LSST and Euclid have the potential to massively increase this sample, with an expected $\sim 1700$ systems observable by the latter. To better understand the significance of the line of sight as a function of the redshifts of the lens and sources, we repeat the procedure described in section 5.1 for a sample of 1729 systems forecasted to be observable with Euclid. For each of these systems, we simulate 1000 lines of sight and calculate the biases on $\eta$ and $g_{\mathrm{LOS}}$ which would arise from ignoring these line-of-sight effects.
@@ 6 Impact on cosmology
Figure 5: Constraints on $w$ and $\Omega_{\mathrm{M}}$ in a flat-$w$CDM cosmology. Shown in blue are the constraints from the Jackpot lens alone, where the uncertainty from the line of sight added in quadrature to the measurement error reported in [15]. In gray are the constraints from Planck [1], and in red are the combined constraints from the Jackpot lens and Planck. The combined constraints give a value of $w=-1.17^{+0.19}_{-0.21}$.
Double-source-plane lenses offer a unique observable in the quantity $\eta$, which allows us to place unique constraints on cosmological quantities such as $\Omega_{\mathrm{M}}$, the matter density parameter, and $w(z)$, the dark energy equation of state [10, 75]. This has motivated the work of [60, 15], and the anticipated samples of DSPLs from LSST and Euclid promise to offer new and competitive constraints [11]. In this section, we aim to explore and quantify the effect of the line of sight on the cosmological constraints possible from DSPLs.
@@ 6.1 Updated constraints from the Jackpot lens
Using $\eta=1.405^{+0.014}_{-0.016}$, $z_{\mathrm{d}}=0.222$, $z_{\mathrm{s_{1}}}=0.609$ and $z_{\mathrm{s_{2}}}=2.41^{+0.04}_{-0.21}$[60] obtained constraints on $\Omega_{\mathrm{M}}$ and $w$ for a flat $w$CDM cosmology, which were updated in [15] with the new spectroscopic redshift measurement for the background source, $z_{\mathrm{s_{2}}}=2.035$. These constraints, however, did not take into account the additional uncertainty on $\eta$ arising from the line of sight. Following the results presented in section 5.1, we add in quadrature a 0.67% error budget for the line of sight to the measurement uncertainty presented in [15], and present the updated constraints in Figure 5. Also shown are the constraints from Planck [1], and the combined constraints from Planck and the Jackpot lens with the line-of-sight error included, giving a dark energy equation of state parameter of $w=-1.17^{+0.19}_{-0.21}$. Using the same data but not including the line-of-sight error, the constraint is instead $w=-1.12^{+0.17}_{-0.18}$, meaning that the line of sight adds $\sim 15\%$ to the uncertainty of the final result.
@@ 6.2 Constraints from a large sample of DSPLs
Figure 6: The scatter introduced by the line of sight compared to the uncertainty in cosmological parameter constraints for a $w$CDM (left) or $w_{0}w_{a}$ (right) cosmology. In green, we show the 1, 2, and 3-$\sigma$ constraints possible from the forecasted set of 1729 Euclid DSPLs and the anticipated measurement uncertainties in $\eta$, with no budget for the line of sight error. In blue, we show the contours for the best guess of these parameters, for 50 000 variations on the lines of sight of each of the DSPLs used in obtaining the constraints. The “true” parameter values are plotted in red.
To properly understand the significance of the line of sight for constraining cosmology with DSPLs, we will consider the forecasted constraints possible from the $\sim$ 1700 such systems expected to be observable with Euclid. If the line of sight is not accounted for in the error budget, but nonetheless present in the data, the main effect is to shift the posterior samples within the parameter space, without significantly affecting their shape or width. If the effect is significantly large, the posterior may be shifted such that the true value no longer lies within 1-$\sigma$ of the measurement. However, given the large number of DSPLs, we would also expect these effects to be mitigated, given that the line of sight bias is not systematic.
To investigate this, we firstly consider the constraining power of the forecasted sample of Euclid systems and the precision with which their $\eta$ values are expected to be measured. This is shown in green in fig. 6 for a flat $w$CDM cosmology (left), as well as for the CPL parameterisation of a redshift-dependent dark energy equation of state,
@@
As expected from fig. 3, the scatter arising from the line of sight is usually subdominant compared to the anticipated measurement uncertainties on the $\eta$ values. In $\sim 95\%$ of line-of-sight realisations, the centroid of the distribution remains within the 1-$\sigma$ errorbars of the idealised constraints for a flat-$w$CDM cosmology. This is also true in the majority of realisations for a flat-$w_{0}w_{a}$ cosmology, although, as shown by the small 2-$\sigma$ region in the lower part of the figure on the right in fig. 6, line-of-sight inhomogeneities are sometimes enough to skew the cosmological constraints more strongly. If line-of-sight effects are neglected, $\sim 35\%$ of realisations of Euclid DSPL lines of sight result in a 2-$\sigma$ tension for $w_{0}$ and $w_{a}$ compared to their true values.
@@ 7 Discussion & conclusion
Finally, we have explored the impact of these biases on the cosmological constraints possible from DSPLs. For a single measurement of $\eta$, the additional uncertainty arising from the line of sight adds a small additional width to the constraints, and we have presented updated flat $w$CDM constraints from J0946+1006, the ‘Jackpot’ lens. Because this additional error is subdominant and expected to be uncorrelated with the measurement error, its impact on these constraints is small. Combining these constraints with those from Planck [1], we obtain a value of $w=-1.17^{+0.19}_{-0.21}$ for the dark energy equation of state parameter, which corresponds to a 14% increase in the reported error when comparing to the case where the line of sight is ignored.
@@
One aspect we have neglected is selection biases. We have assumed that the discovery probablility of a compound lens does not depend significantly on the properties of the line of sight. This is not true for lensed point sources [78, 79], but is less likely to play a role for lensed extended sources [73].
@@
Taking these two actions will ensure that double source plane lens systems become a competitive method for constraining the dark energy equation of state. Of course, our work has only focused on line-of-sight lensing: efforts are still needed to constrain the density profiles of lens galaxies, ensuring that no internal mass-sheet transformation has been missed during lens modelling.
@@ Acknowledgements
DJ acknowledges support by the First Rand Foundation, South Africa, and the Centre National de la Recherche Scientifique of France. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LensEra: grant agreement No. 945536). TC is funded by the Royal Society through a University Research Fellowship. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising. PF acknowledges support from the French Agence Nationale de la Recherche through the ELROND project (ANR-23-CE31-0002).
@@
Daniel Johnson: Methodology, Formal Analysis, Writing - Original draft. Thomas Collett: Conceptualization, Methodology, Writing - Review & Editing, Supervision. Tian Li: Software Pierre Fleury: Methodology, Supervision

### v2

@@ 1 Introduction
[…]
@@
The first observation of such a system, The ‘Jackpot’ lens J0946+1006, was identified by [12] with the Sloan Digital Sky Survey [13] and follow ups with the Hubble Space Telescope. Later, a third, more distant multiply-imaged source was observed as part of the same system [14, 15], making J0946+1006 also the first triple source system. This system has been used to place constraints on $\Omega_{\mathrm{m}}$ and the dark energy equation of state parameter $w$. Most recently, re-analysis of the known galaxy-scale strong lensing system J1721+8842 revealed that the six observed quasar images arose from the same background source, the first confirmation of an Einstein zig-zag lens [16].
While only five other DSPL systems have been spectroscopically confirmed [17, 18, 19, 20, 21], upcoming surveys are expected to increase this number by at least two orders of magnitude [11]. The Legacy Survey of Space and Time (LSST) [22] and the Euclid [23] telescope are both forecast to substantially increase the strong lens population [24]. Euclid is particularly powerful for DSPL discovery, since it will survey roughly a third of the entire sky with a 0.1 arcsecond resolution, unprecedented for such a wide survey. Identifying a DSPL likely requires the Einstein radii of each source to differ by more than the angular resolution, so Euclid’s excellent resolution offers a major advantage for DSPL identification. Euclid is forecast to detect $\sim$ 1700 such systems [25], and 4 have been serendipitously discovered in the initial Q1 area [25] (which is a representative patch of 64 deg2 of the final $\sim$140000 deg2 Euclid wide survey [26]) . Given the high number of systems, the extensive sky coverage and the possibility of precise redshift measurements of lens and source light, Euclid holds massive promise for DSPL cosmology [11].
In recent years, increasing attention has been paid to the impact on strong lensing observables of inhomogeneities along the line of sight, and it is now understood that, for precision applications of strong lensing, these effects cannot be ignored [27, 28, 29, 30, 31]. Gravitational shear effects arising from matter external to the main lens [32, 33, 34, 35, 36, 37] can be as strong as the shear produced by the main lens itself [37], and a few percent of the total lensing distortion [38]. Consequently, azimuthal weak lensing effects must be included in models of strong lenses, and their omission can introduce biases [39, 40, 41] or prevent successful modelling [39]. The most common approach is to include a tidal “external shear” term, meant to absorb these effects, as well as any shear not reproduced by the mass model of the main lens [42, 43, 44, 37].
@@
This paper is organised as follows: in section 2, we review the theory of double source plane lensing and its use in constraining cosmology, and show how line-of-sight inhomogeneities affect DSPL observables. In section 3, we present our methods for simulating realistic tidal line of sight effects for DSPLs with arbitrary redshifts, and in section 4, we present the sample of DSPLs expected to be observed with Euclid. In section 5, we quantify the impact of the line of sight on the cosmological scaling parameter and the line of sight shear observed from DSPL systems. In section 6, we present flat-wwCDM cosmological constraints from the Jackpot lens with an error budget updated to include a contribution from the line of sight, and then explore the impact of the line of sight on flat-wwCDM and $w_{0}w_{a}$CDM cosmological constraints from the forecasted Euclid DSPLs. We summarise our results in section 7.
@@ 2.1.2 Cosmology with DSPLs
For a wwCDM cosmology (with $w$ constant but not necessarily equal to 1-1), the normalised Hubble parameter $E(z)\equiv H(z)/H_{0}$ is given by
@@ 2.3.1 The cosmological scaling factor
where the $\sigma_{v}^{2}$ are the velocity dispersions of d and s1. Clearly, this inequality need not be satisfied in general.
@@ 4 Forecasted double source plane lens populations
Euclid is expected to contain a sample of $\sim 1700$ galaxy-scale DSPL systems that could plausibly be used to constrain cosmological parameters [25]. This forecast was derived from the LensPop package [24] modified to include multiple background sources and neglecting mass in the first source. This sample only includes systems where both sources have one or more arcs of length 0.3, which ensures a reasonable possibility that the density slope of the lens can be recovered from Euclid imaging alone. In this work, we will assume that Euclid is able to find all of these forecasted lenses, with redshifts and Einstein radii following [25]. As in [25], we will assume that each DSPL enables a measurement of $\eta$ with an uncertainty given by the quadrature sum of 0.01 arcseconds over the Einstein radius of the first source, the same for the second source and 0.01. The first two terms represent reasonable uncertainties on the Einstein radius determined from lens modelling, whilst the last term sets a floor on the uncertainty given that the density profile slope of the primary lens is not perfectly known.
@@ 5.1 A single system - the Jackpot lens
From the figure on the left, we see that the line of sight typically introduces a bias on $\eta$ on the order of a percent, with a standard deviation of $\eta^{\mathrm{inferred}}/\eta$ of 0.67% for the inner pair of rings, and 1.3% for the outer. While the median of the bias is $\approx 0$, the histograms are not perfectly symmetric, and large overestimations of $\eta$ are more frequent than large underestimations. Unsurprisingly, we see that the typical size of the bias increases when considering the value of $\eta$ as calculated between s1 and s3, as the larger redshift differences between these two sources mean the line of sight becomes, on average, more significant (see fig. 1). For comparison, the measurement error in $\eta$ presented in [65] is 1.1%. That statistical error thus remains the dominant source of uncertainty, but the effect of the line of sight is not completely negligible. While our lines of sight are chosen without accounting for selection effects, some simple tests suggest that such effects may add a small additional bias to measurements of $\eta$, but that this bias is unlikely to be much larger than the scatter seen in fig. 2. This is discussed in more detail in appendix B.
From the figure on the right, we see that the difference between the line-of-sight shear acting on the inner and outer rings is typically slightly smaller than a percent, once again growing larger when comparing with the more distant source. While these values are small in absolute terms, this corresponds to a median difference of $\sim$140% of $\eta\,g_{\mathrm{LOS,1}}$ in the case of s2, and $\sim$230% in the case of s3. It is therefore clear that $\eta\,g_{\mathrm{LOS,1}}$ is a poor substitute when modelling $g_{\mathrm{LOS,2}}$ and $g_{\mathrm{LOS,3}}$. Compound lens models should allow for an additional external shear term on each lens plane.
@@ 5.2 A sample of mock Euclid DSPLs
To better understand the significance of the line of sight as a function of the redshifts of the lens and sources, we repeat the procedure described in section 5.1 for a sample of 1729 systems forecasted to be observable with Euclid (section 4). For each of these systems, we simulate 1000 lines of sight and calculate the biases on $\eta$ and $g_{\mathrm{LOS}}$ which would arise from ignoring these line-of-sight effects.
@@ 6 Impact on cosmology
Figure 5: Constraints on $w$ and $\Omega_{\mathrm{m}}$ in a flat-wwCDM cosmology. Shown in blue are the constraints from the Jackpot lens alone, where the uncertainty from the line of sight added in quadrature to the measurement error reported in [15]. In gray are the constraints from Planck [1], and in red are the combined constraints from the Jackpot lens and Planck. The combined constraints give a value of $w=-1.17^{+0.19}_{-0.21}$.
Measurements of the quantity $\eta$ from DSPLs allow us to place constraints on cosmological quantities such as $\Omega_{\mathrm{m}}$, the matter density parameter, and $w(z)$, the dark energy equation of state [10, 78]. This has motivated the work of [65, 15], and the anticipated samples of DSPLs from LSST and Euclid promise to offer new and competitive constraints [11]. In this section, we aim to explore and quantify the effect of the line of sight on the cosmological constraints possible from DSPLs.
@@ 6.1 Updated constraints from the Jackpot lens
Using $\eta=1.405^{+0.014}_{-0.016}$, $z_{\mathrm{d}}=0.222$, $z_{\mathrm{s_{1}}}=0.609$ and $z_{\mathrm{s_{2}}}=2.41^{+0.04}_{-0.21}$[65] obtained constraints on $\Omega_{\mathrm{m}}$ and $w$ for a flat wwCDM cosmology, which were updated in [15] with the new spectroscopic redshift measurement for the background source, $z_{\mathrm{s_{2}}}=2.035$. These constraints, however, did not take into account the additional uncertainty on $\eta$ arising from the line of sight. Following the results presented in section 5.1, we add in quadrature a 0.67% error budget for the line of sight to the measurement uncertainty presented in [15], and present the updated constraints in Figure 5. Also shown are the constraints from Planck [1], and the combined constraints from Planck and the Jackpot lens with the line-of-sight error included, giving a dark energy equation of state parameter of $w=-1.17^{+0.19}_{-0.21}$. Using the same data but not including the line-of-sight error, the constraint is instead $w=-1.12^{+0.17}_{-0.18}$, meaning that the line of sight adds $\sim 15\%$ to the uncertainty of the final result.
@@ 6.2 Constraints from a large sample of DSPLs
Figure 6: The scatter introduced by the line of sight compared to the uncertainty in cosmological parameter constraints for a wwCDM (left) or $w_{0}w_{a}$CDM (right) cosmology. In green, we show the 1, 2, and 3-$\sigma$ constraints possible from the forecasted set of 1729 Euclid DSPLs and the anticipated measurement uncertainties in $\eta$, with no budget for the line of sight error. In blue, we show the contours for the best guess of these parameters, for 50 000 variations on the lines of sight of each of the DSPLs used in obtaining the constraints. The “true” parameter values are plotted in red.
We now quantify the impact of these line-of-sight uncertainties on the cosmological parameter constraints possible from a Euclid-like DSPL sample. If the line of sight is not accounted for in the error budget for $\eta$, but nonetheless present in the data, the main effect is to shift the posterior samples within the parameter space, without significantly affecting their shape or width. If the effect is significantly large, the posterior may be shifted such that the true value no longer lies within 1-$\sigma$ of the measurement. However, given the large number of DSPLs, we would also expect these effects to be mitigated, given that the line of sight bias is not systematic.
To investigate this, we firstly consider the constraining power of the forecasted sample of Euclid systems and the precision with which their $\eta$ values are expected to be measured. This is shown in green in fig. 6 for a flat wwCDM cosmology (left), as well as for the CPL parameterisation of a redshift-dependent dark energy equation of state,
@@
As expected from fig. 3, the scatter arising from the line of sight is usually subdominant compared to the anticipated measurement uncertainties on the $\eta$ values. In $\sim 95\%$ of line-of-sight realisations, the centroid of the distribution remains within the 1-$\sigma$ errorbars of the idealised constraints for a flat-wwCDM cosmology. This is also true in the majority of realisations for a flat-$w_{0}w_{a}$CDM cosmology, although, as shown by the small 2-$\sigma$ region in the lower part of the figure on the right in fig. 6, line-of-sight inhomogeneities are sometimes enough to skew the cosmological constraints more strongly. If line-of-sight effects are neglected, $\sim 35\%$ of realisations of Euclid DSPL lines of sight result in a 2-$\sigma$ tension for $w_{0}$ and $w_{a}$ compared to their true values.
It is beyond the scope of this work to fully account for the selection effects which may bias the distribution of lines of sight. Nonetheless, some simple considerations show that, while the number of outliers may be increased when accounting for the effect of line-of-sight influenced Einstein radii on the probability of DSPLs occurring, this does not seem to introduce a noticable systematic bias in cosmological constraints. This is discussed and illustrated in appendix B.
@@ 7 Discussion & conclusion
Finally, we have explored the impact of these biases on the cosmological constraints possible from DSPLs. For a single measurement of $\eta$, the additional uncertainty arising from the line of sight adds a small additional width to the constraints, and we have presented updated flat wwCDM constraints from J0946+1006, the ‘Jackpot’ lens. Because this additional error is subdominant and expected to be uncorrelated with the measurement error, its impact on these constraints is small. Combining these constraints with those from Planck [1], we obtain a value of $w=-1.17^{+0.19}_{-0.21}$ for the dark energy equation of state parameter, which corresponds to a 14% increase in the reported error when comparing to the case where the line of sight is ignored.
@@
Our investigation of links between the properties of the line of sight to the occurrence and discovery of DSPLs has been limited to some simple illustrative models, and is mostly presented in appendix B. Work remains to be done to thoroughly quantify selection effects at play in double-source-plane lensing, and the tools necessary for this are beyond the scope of this work. While the lensing of point sources is known to be sensitive to these effects [81, 82], they are understood to play a smaller role for lensed extended sources [24]. This is consistent with our findings, that the biases introduced tend to be subdominant when compared to the uncertainty arising from the line of sight.
While our work has focused on line-of-sight lensing, efforts are still needed to constrain the density profiles of lens galaxies, ensuring that no internal mass-sheet transformation has been missed during lens modelling. A mass sheet in the lens plane need not correspond in any way to expectations from the distributions of lines of sight on larger scales, and, like the line-of-sight effects discussed in this work, can lead to precise but inaccurate lens reconstructions and cosmological constraints [62].
@@
Taking these two actions will ensure that double source plane lens systems become a competitive method for constraining the dark energy equation of state.
@@ Acknowledgements
DJ acknowledges support by the First Rand Foundation, South Africa, and the Centre National de la Recherche Scientifique of France. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LensEra: grant agreement No. 945536). TC is funded by the Royal Society through a University Research Fellowship. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising. PF acknowledges support from the French Agence Nationale de la Recherche through the ELROND project (ANR-23-CE31-0002). We thank the anonymous reviewer for their helpful comments on an earlier version of this work.
@@
Daniel Johnson: Methodology, Formal Analysis, Writing - Original draft. Thomas Collett: Conceptualization, Methodology, Writing - Review & Editing, Supervision. Tian Li: Software. Pierre Fleury: Methodology, Supervision.